# Deep Modularity Networks with Diversity–Preserving Regularization

### v1

@@ A.1 Varying the Epsilon
A.2 Varying the Variance Weight
Figure 4 depicts the ablation study done to find the best value of $W_{\text{var}}$ by setting $W_{\text{dist}}$ and $W_{\text{entropy}}$ weights to 0. For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $W_{\text{var}}$ values were found to be $1$/$0.1$, $0.1$, $0.001$, $0.1$, and $1.0$ respectively.
@@ A.2 Varying the Variance Weight
A.3 Varying the Entropy Weight
@@
Appendix B Full Results
The tables below are indicative of the value of each evaluation metric obtained using DMoN/DMoN–DPR pooling for each individual seed for different datasets. The numbers are all in percentage (%).
B.1 Cora
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Cora dataset.
@@ B.1 Cora
993 DMoN 10.25 73.93 48.71 55.38
550 DMoN 9.95 71.77 40.62 42.17
243 DMoN 10.06 73.74 47.62 53.20
16 DMoN 12.13 70.62 44.77 46.61
716 DMoN 11.92 70.40 37.87 39.65
383 DMoN 11.20 71.80 44.40 52.46
277 DMoN 11.03 72.50 37.07 36.07
274 DMoN 10.12 74.14 45.30 44.74
188 DMoN 11.48 73.36 46.79 49.47
796 DMoN 11.86 71.67 43.30 42.69
Mean 11.00 72.39 43.65 46.24
Variance 0.01 0.02 0.16 0.40
@@
993 DMoN–DPR (D) 10.50 73.78 51.01 57.45
550 DMoN–DPR (D) 10.06 72.20 41.49 44.43
243 DMoN–DPR (D) 9.49 73.58 47.80 54.94
16 DMoN–DPR (D) 10.55 72.20 46.22 49.19
716 DMoN–DPR (D) 10.38 71.10 39.29 43.43
383 DMoN–DPR (D) 9.49 74.11 48.02 54.77
277 DMoN–DPR (D) 10.65 72.45 40.12 42.77
274 DMoN–DPR (D) 9.47 74.41 47.64 52.47
188 DMoN–DPR (D) 10.44 73.75 48.41 54.78
796 DMoN–DPR (D) 11.29 72.13 43.05 45.75
Mean 10.23 72.97 45.31 50.00
Variance 0.00 0.01 0.16 0.31
@@
993 DMoN–DPR 10.99 72.41 51.86 58.92
550 DMoN–DPR (V) 11.10 70.31 40.04 43.09
243 DMoN–DPR (V) 11.75 71.02 47.47 54.54
16 DMoN–DPR (V) 10.99 70.18 45.82 48.81
716 DMoN–DPR (V) 11.08 69.84 40.82 41.88
383 DMoN–DPR (V) 12.13 69.37 42.89 51.12
277 DMoN–DPR (V) 11.42 70.91 37.43 37.64
274 DMoN–DPR (V) 11.05 72.61 46.97 48.40
188 DMoN–DPR (V) 13.11 70.74 45.71 52.26
796 DMoN–DPR (V) 12.79 70.56 43.70 45.13
Mean 11.64 70.80 44.27 48.18
Variance 0.01 0.01 0.18 0.41
@@
993 DMoN–DPR (E) 10.25 73.81 49.10 55.86
550 DMoN–DPR (E) 10.06 71.70 40.49 42.06
243 DMoN–DPR (E) 9.91 73.82 48.60 53.72
16 DMoN–DPR (E) 12.32 70.46 44.95 46.88
716 DMoN–DPR (E) 11.86 70.43 37.72 39.66
383 DMoN–DPR (E) 11.01 71.93 44.49 52.83
277 DMoN–DPR (E) 11.16 72.37 37.38 35.98
274 DMoN–DPR (E) 10.17 74.20 44.13 44.02
188 DMoN–DPR (E) 11.41 73.37 46.28 48.86
796 DMoN–DPR (E) 11.88 71.65 43.17 42.63
Mean 11.00 72.37 43.63 46.25
Variance 0.01 0.02 0.16 0.43
@@
993 DMoN–DPR (DV) 10.25 73.42 51.71 58.37
550 DMoN–DPR (DV) 10.17 72.08 41.28 44.08
243 DMoN–DPR (DV) 9.66 73.18 48.08 55.94
16 DMoN–DPR (DV) 10.72 71.69 46.80 49.51
716 DMoN–DPR (DV) 10.52 70.97 39.46 43.32
383 DMoN–DPR (DV) 9.87 73.82 48.11 54.87
277 DMoN–DPR (DV) 10.95 71.90 39.01 42.43
274 DMoN–DPR (DV) 9.61 74.13 48.04 53.26
188 DMoN–DPR (DV) 10.57 73.51 48.47 54.55
796 DMoN–DPR (DV) 11.39 72.06 43.12 44.93
Mean 10.37 72.68 45.41 50.13
Variance 0.00 0.01 0.19 0.36
@@
993 DMoN–DPR (DE) 10.63 73.82 51.30 57.10
550 DMoN–DPR (DE) 10.10 72.15 41.06 44.04
243 DMoN–DPR (DE) 9.68 73.25 47.62 55.30
16 DMoN–DPR (DE) 10.61 72.13 46.23 49.09
716 DMoN–DPR (DE) 10.63 70.90 39.37 43.28
383 DMoN–DPR (DE) 9.53 74.04 47.93 54.43
277 DMoN–DPR (DE) 10.89 72.25 39.14 42.53
274 DMoN–DPR (DE) 9.49 74.39 48.31 52.84
188 DMoN–DPR (DE) 10.52 73.68 48.16 54.61
796 DMoN–DPR (DE) 11.18 72.21 42.71 45.56
Mean 10.33 72.88 45.18 49.88
Variance 0.00 0.01 0.18 0.32
@@
993 DMoN–DPR (VE) 10.86 72.34 51.43 58.35
550 DMoN–DPR (VE) 11.10 70.30 40.10 43.20
243 DMoN–DPR (VE) 11.71 71.05 47.46 54.42
16 DMoN–DPR (VE) 11.06 70.19 45.89 48.85
716 DMoN–DPR (VE) 11.01 69.83 40.77 41.82
383 DMoN–DPR (VE) 12.14 69.28 42.65 50.92
277 DMoN–DPR (VE) 11.46 70.98 37.40 37.53
274 DMoN–DPR (VE) 11.14 72.47 46.72 48.43
188 DMoN–DPR (VE) 13.11 70.78 45.36 52.19
796 DMoN–DPR (VE) 12.85 70.48 43.74 45.16
Mean 11.64 70.77 44.15 48.09
Variance 0.01 0.01 0.17 0.39
@@
993 DMoN–DPR (DVE) 10.06 73.55 51.62 58.48
550 DMoN–DPR (DVE) 10.16 72.04 41.47 44.30
243 DMoN–DPR (DVE) 9.87 72.96 47.86 55.57
16 DMoN–DPR (DVE) 10.76 71.64 46.63 49.25
716 DMoN–DPR (DVE) 10.72 70.81 39.64 43.32
383 DMoN–DPR (DVE) 9.93 73.75 48.14 54.86
277 DMoN–DPR (DVE) 10.89 71.99 39.10 42.50
274 DMoN–DPR (DVE) 9.83 73.94 47.81 52.85
188 DMoN–DPR (DVE) 10.42 73.73 48.53 54.50
796 DMoN–DPR (DVE) 11.41 72.05 43.51 44.96
Mean 10.41 72.65 45.43 50.06
Variance 0.00 0.01 0.18 0.35
B.2 CiteSeer
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the CiteSeer dataset.
@@ B.2 CiteSeer
993 DMoN 6.09 74.10 22.56 26.60
550 DMoN 4.31 75.22 25.92 41.43
243 DMoN 4.90 78.26 34.20 46.62
16 DMoN 5.87 75.48 28.89 38.28
716 DMoN 3.60 76.57 38.40 52.01
383 DMoN 4.94 75.34 26.50 39.64
277 DMoN 6.50 75.67 28.87 37.96
274 DMoN 4.35 74.93 32.36 47.53
188 DMoN 4.50 76.41 34.86 46.98
796 DMoN 4.13 75.24 30.07 44.67
Mean 4.92 75.72 30.26 42.17
Variance 0.01 0.01 0.23 0.51
@@
993 DMoN–DPR (D) 6.30 74.07 22.96 27.06
550 DMoN–DPR (D) 4.11 75.45 26.52 42.31
243 DMoN–DPR (D) 3.98 76.17 36.63 51.64
16 DMoN–DPR (D) 5.36 76.07 29.32 42.93
716 DMoN–DPR (D) 4.17 76.07 39.05 51.65
383 DMoN–DPR (D) 5.18 75.60 26.83 37.54
277 DMoN–DPR (D) 5.16 75.30 29.25 40.17
274 DMoN–DPR (D) 4.28 75.19 32.02 46.99
188 DMoN–DPR (D) 4.70 76.60 34.34 45.97
796 DMoN–DPR (D) 4.24 75.92 29.15 41.29
Mean 4.75 75.64 30.61 42.76
Variance 0.01 0.00 0.24 0.52
@@
993 DMoN–DPR (V) 6.33 73.98 23.33 26.50
550 DMoN–DPR (V) 4.17 75.20 26.64 42.18
243 DMoN–DPR (V) 4.99 78.15 34.08 46.34
16 DMoN–DPR (V) 6.02 75.21 29.38 38.80
716 DMoN–DPR (V) 3.71 76.45 38.70 52.38
383 DMoN–DPR (V) 5.16 74.60 26.80 39.65
277 DMoN–DPR (V) 5.84 75.26 29.13 39.42
274 DMoN–DPR (V) 4.42 74.79 31.87 46.92
188 DMoN–DPR (V) 4.33 76.39 34.76 47.14
796 DMoN–DPR (V) 4.33 75.87 29.92 43.30
Mean 4.93 75.59 30.46 42.26
Variance 0.01 0.01 0.20 0.49
@@
993 DMoN–DPR (E) 6.59 74.00 22.17 27.07
550 DMoN–DPR (E) 4.42 75.06 26.32 41.61
243 DMoN–DPR (E) 5.14 77.95 34.48 46.79
16 DMoN–DPR (E) 6.15 75.09 29.46 38.54
716 DMoN–DPR (E) 3.67 76.36 39.12 52.80
383 DMoN–DPR (E) 5.32 74.52 27.05 39.70
277 DMoN–DPR (E) 5.89 75.27 29.14 39.34
274 DMoN–DPR (E) 4.35 74.85 32.18 47.28
188 DMoN–DPR (E) 4.55 76.16 34.69 47.23
796 DMoN–DPR (E) 4.11 75.02 30.58 45.06
Mean 5.02 75.43 30.52 42.54
Variance 0.01 0.01 0.24 0.50
@@
993 DMoN–DPR (DV) 6.33 73.81 23.64 27.24
550 DMoN–DPR (DV) 4.09 75.48 26.65 42.08
243 DMoN–DPR (DV) 3.89 76.17 35.67 51.00
16 DMoN–DPR (DV) 5.71 75.73 29.33 42.91
716 DMoN–DPR (DV) 4.11 76.03 38.91 51.73
383 DMoN–DPR (DV) 5.16 75.31 27.06 38.22
277 DMoN–DPR (DV) 5.01 75.32 29.81 40.68
274 DMoN–DPR (DV) 4.35 74.92 32.29 47.22
188 DMoN–DPR (DV) 4.50 76.36 34.12 46.47
796 DMoN–DPR (DV) 4.44 75.93 28.46 40.94
Mean 4.76 75.51 30.59 42.85
Variance 0.01 0.01 0.21 0.50
@@
993 DMoN–DPR (DE) 6.48 73.71 23.56 27.55
550 DMoN–DPR (DE) 4.31 75.24 26.29 41.96
243 DMoN–DPR (DE) 3.95 76.11 36.10 51.46
16 DMoN–DPR (DE) 5.76 75.68 29.78 43.37
716 DMoN–DPR (DE) 4.15 76.00 39.03 51.88
383 DMoN–DPR (DE) 5.43 75.09 27.51 38.72
277 DMoN–DPR (DE) 5.10 75.25 29.91 40.77
274 DMoN–DPR (DE) 4.42 74.89 31.85 46.74
188 DMoN–DPR (DE) 4.75 76.28 34.05 46.33
796 DMoN–DPR (DE) 4.26 75.21 29.63 43.94
Mean 4.86 75.35 30.77 43.27
Variance 0.01 0.01 0.22 0.49
@@
993 DMoN–DPR (VE) 6.50 73.55 22.91 27.80
550 DMoN–DPR (VE) 4.59 74.88 26.41 41.35
243 DMoN–DPR (VE) 5.25 77.84 34.60 47.05
16 DMoN–DPR (VE) 5.95 76.05 31.65 40.71
716 DMoN–DPR (VE) 3.84 76.18 39.28 52.93
383 DMoN–DPR (VE) 5.40 74.12 26.44 39.40
277 DMoN–DPR (VE) 5.91 75.11 29.14 39.27
274 DMoN–DPR (VE) 4.55 74.52 32.02 47.12
188 DMoN–DPR (VE) 4.70 75.85 34.92 47.37
796 DMoN–DPR (VE) 4.24 74.50 31.26 45.61
Mean 5.09 75.26 30.86 42.86
Variance 0.01 0.02 0.23 0.47
@@
993 DMoN–DPR (DVE) 6.39 73.36 23.60 28.45
550 DMoN–DPR (DVE) 4.35 75.08 26.35 41.77
243 DMoN–DPR (DVE) 3.93 75.80 36.58 52.14
16 DMoN–DPR (DVE) 5.38 76.03 28.50 41.56
716 DMoN–DPR (DVE) 4.22 75.83 38.91 51.92
383 DMoN–DPR (DVE) 5.34 74.92 28.02 39.07
277 DMoN–DPR (DVE) 5.34 75.06 29.87 40.58
274 DMoN–DPR (DVE) 4.46 74.65 31.76 46.98
188 DMoN–DPR (DVE) 4.68 75.90 34.46 47.48
796 DMoN–DPR (DVE) 4.44 74.79 29.69 44.27
Mean 4.85 75.14 30.77 43.42
Variance 0.01 0.01 0.22 0.48
B.3 PubMed
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the PubMed dataset.

### v2

@@ D.1.1 Varying the Epsilon
D.1.2 Varying the Variance Weight
Figure 3 depicts the ablation study done to find the best value of $W_{\text{var}}$ by setting $W_{\text{dist}}$ and $W_{\text{entropy}}$ weights to 0. For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $W_{\text{var}}$ values were found to be 11/$0.1$, $0.1$, $0.001$, $0.1$, and $1.0$ respectively.
@@ D.1.2 Varying the Variance Weight
D.1.3 Varying the Entropy Weight
@@
D.1.4 Practical Challenges Regarding Manual Tuning
We acknowledge that adding diversity‐preserving regularizers introduces more hyperparameters, which can complicate practical adoption. However, in our experiments:
  • We found that larger coefficients for distance (R_D = 1 or 10) were typically beneficial on highly feature‐rich datasets (Coauthor CS, Coauthor Physics).
  • Coefficients of 1 to 0.1 for variance (R_V) often worked well across most datasets.
  • For the entropy term, 0.1 was a good upper bound on large feature spaces, whereas a smaller coefficient was better on less feature‐rich datasets.
From these observations, we hypothesize that stronger diversity coefficients are advantageous for richer feature spaces, whereas smaller coefficients are sufficient (and sometimes necessary) for lower‐dimensional or simpler datasets. As a direction for future work, we suggest exploring automated or adaptive tuning schemes. Such an approach could dynamically adjust the weight of each regularizer during training, further reducing manual overhead and improving reproducibility.
Appendix E Full Results
The tables below are indicative of the value of each evaluation metric obtained using DMoN/DMoN–DPR pooling for each individual seed for different datasets. The seeds were selected at random using a random number generator. The numbers are all in percentage (%).
E.1 Cora
The following tables list the performance of DiffPool, MinCut Pool, DMoN and different DMoN–DPR configurations across different seeds on the Cora dataset.
Table 8: Results obtained by using DiffPool on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DiffPool 15.90 63.81 42.06 51.72
550 DiffPool 18.74 61.16 35.22 36.23
243 DiffPool 15.84 64.02 40.57 49.64
16 DiffPool 17.07 57.90 37.43 46.73
716 DiffPool 15.33 64.34 46.69 53.05
383 DiffPool 15.54 63.93 39.65 44.05
277 DiffPool 12.83 67.00 47.31 55.25
274 DiffPool 16.31 63.31 36.01 39.97
188 DiffPool 18.26 56.66 31.67 35.33
796 DiffPool 14.08 65.66 44.65 53.50
Mean $\pm$ Std 15.99 $\pm$ 1.77 62.78 $\pm$ 3.28 40.13 $\pm$ 5.15 46.55 $\pm$ 7.34
Table 9: Results obtained by using MinCut pooling on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 MinCut 12.73 72.59 41.66 43.90
550 MinCut 11.61 73.70 45.63 45.03
243 MinCut 13.89 71.56 31.48 33.71
16 MinCut 14.08 71.37 38.06 37.21
716 MinCut 14.38 71.14 33.82 40.37
383 MinCut 15.18 70.36 29.04 28.48
277 MinCut 13.09 72.42 42.81 46.15
274 MinCut 13.32 72.16 39.60 39.77
188 MinCut 13.81 71.65 40.52 44.13
796 MinCut 14.44 70.91 34.75 32.71
Mean $\pm$ Std 13.65 $\pm$ 1.01 71.79 $\pm$ 0.96 37.74 $\pm$ 5.31 39.15 $\pm$ 5.98
@@ E.1 Cora
993 DMoN–DPR (DVE) 11.35 71.32 48.90 52.56
550 DMoN–DPR (DVE) 9.81 72.13 38.15 42.84
243 DMoN–DPR (DVE) 10.72 73.08 40.71 40.03
16 DMoN–DPR (DVE) 11.06 71.48 42.84 46.41
716 DMoN–DPR (DVE) 9.93 72.55 42.77 44.43
383 DMoN–DPR (DVE) 11.54 73.41 44.12 51.88
277 DMoN–DPR (DVE) 9.66 73.52 45.61 49.43
274 DMoN–DPR (DVE) 11.03 74.44 41.98 41.83
188 DMoN–DPR (DVE) 9.51 74.37 51.31 56.52
796 DMoN–DPR (DVE) 10.65 71.36 42.78 43.37
Mean $\pm$ Std 10.53 $\pm$ 0.74 72.77 $\pm$ 1.18 43.92 $\pm$ 3.85 46.93 $\pm$ 5.42
@@
993 DMoN–DPR (D) 11.31 71.37 49.62 53.10
550 DMoN–DPR (D) 9.76 72.28 40.96 45.72
243 DMoN–DPR (D) 10.40 74.23 43.14 44.89
16 DMoN–DPR (D) 11.90 71.00 42.22 44.14
716 DMoN–DPR (D) 9.62 71.51 43.06 46.72
383 DMoN–DPR (D) 11.24 72.86 43.27 51.09
277 DMoN–DPR (D) 13.64 69.81 40.51 44.53
274 DMoN–DPR (D) 11.12 73.24 43.58 45.78
188 DMoN–DPR (D) 9.80 74.15 50.89 56.19
796 DMoN–DPR (D) 10.70 71.52 42.57 42.95
Mean $\pm$ Std 10.95 $\pm$ 1.21 72.20 $\pm$ 1.42 43.98 $\pm$ 3.46 47.51 $\pm$ 4.40
@@
993 DMoN–DPR (V) 11.25 70.12 48.19 52.60
550 DMoN–DPR (V) 10.95 70.66 39.21 43.64
243 DMoN–DPR (V) 11.63 71.41 40.33 39.36
16 DMoN–DPR (V) 11.25 70.02 43.27 45.92
716 DMoN–DPR (V) 11.22 70.58 43.18 45.43
383 DMoN–DPR (V) 13.40 70.06 42.14 51.60
277 DMoN–DPR (V) 11.33 69.87 41.46 45.44
274 DMoN–DPR (V) 12.43 72.57 45.58 44.85
188 DMoN–DPR (V) 12.83 70.46 46.59 51.89
796 DMoN–DPR (V) 12.37 69.61 43.49 43.14
Mean $\pm$ Std 11.87 $\pm$ 0.83 70.54 $\pm$ 0.88 43.34 $\pm$ 2.79 46.39 $\pm$ 4.32
@@
993 DMoN–DPR (E) 11.20 71.38 48.47 52.36
550 DMoN–DPR (E) 9.91 72.05 38.08 42.93
243 DMoN–DPR (E) 10.57 73.23 40.85 39.97
16 DMoN–DPR (E) 11.06 71.47 42.93 46.56
716 DMoN–DPR (E) 9.89 72.54 42.62 44.72
383 DMoN–DPR (E) 11.42 73.51 44.30 51.74
277 DMoN–DPR (E) 9.55 73.52 45.49 49.53
274 DMoN–DPR (E) 11.08 74.33 43.10 42.42
188 DMoN–DPR (E) 9.61 74.26 51.24 56.67
796 DMoN–DPR (E) 10.63 71.45 42.93 43.32
Mean $\pm$ Std 10.49 $\pm$ 0.70 72.77 $\pm$ 1.15 44.00 $\pm$ 3.72 47.02 $\pm$ 5.34
@@
993 DMoN–DPR (DV) 10.93 71.05 49.87 53.66
550 DMoN–DPR (DV) 9.51 72.33 41.04 45.65
243 DMoN–DPR (DV) 10.84 73.28 42.18 42.58
16 DMoN–DPR (DV) 10.86 71.23 43.08 44.77
716 DMoN–DPR (DV) 9.61 71.39 44.18 47.59
383 DMoN–DPR (DV) 11.80 73.16 43.66 46.75
277 DMoN–DPR (DV) 13.74 69.61 40.23 44.66
274 DMoN–DPR (DV) 11.33 73.29 45.41 47.47
188 DMoN–DPR (DV) 9.61 74.03 51.32 56.96
796 DMoN–DPR (DV) 10.72 71.30 43.02 43.54
Mean $\pm$ Std 10.89 $\pm$ 1.26 72.07 $\pm$ 1.37 44.40 $\pm$ 3.60 47.36 $\pm$ 4.55
@@
993 DMoN–DPR (DE) 11.16 71.38 49.75 53.29
550 DMoN–DPR (DE) 9.68 72.33 40.85 45.67
243 DMoN–DPR (DE) 10.08 74.45 43.94 46.11
16 DMoN–DPR (DE) 11.29 71.57 43.55 43.97
716 DMoN–DPR (DE) 9.61 71.58 44.19 47.71
383 DMoN–DPR (DE) 11.73 73.06 42.66 48.52
277 DMoN–DPR (DE) 13.58 69.91 40.43 44.48
274 DMoN–DPR (DE) 11.01 73.29 43.61 45.77
188 DMoN–DPR (DE) 9.81 74.04 51.34 56.67
796 DMoN–DPR (DE) 10.82 71.36 42.67 43.10
Mean $\pm$ Std 10.88 $\pm$ 1.21 72.30 $\pm$ 1.40 44.30 $\pm$ 3.54 47.53 $\pm$ 4.32
@@
993 DMoN–DPR (VE) 11.16 70.08 48.03 52.85
550 DMoN–DPR (VE) 10.86 70.77 39.39 43.76
243 DMoN–DPR (VE) 11.78 71.27 40.17 39.23
16 DMoN–DPR (VE) 11.29 69.93 43.33 45.86
716 DMoN–DPR (VE) 11.10 70.70 43.42 45.66
383 DMoN–DPR (VE) 13.17 70.14 41.54 51.22
277 DMoN–DPR (VE) 11.46 69.60 41.28 45.35
274 DMoN–DPR (VE) 12.43 72.56 45.66 44.94
188 DMoN–DPR (VE) 12.90 70.39 46.46 51.76
796 DMoN–DPR (VE) 12.41 69.61 43.10 42.98
Mean $\pm$ Std 11.86 $\pm$ 0.82 70.50 $\pm$ 0.89 43.24 $\pm$ 2.79 46.36 $\pm$ 4.32
@@
993 DMoN–DPR (DVE) 10.65 71.20 49.95 53.90
550 DMoN–DPR (DVE) 9.51 72.33 41.04 45.64
243 DMoN–DPR (DVE) 10.86 73.34 41.92 42.43
16 DMoN–DPR (DVE) 10.78 71.29 42.83 44.73
716 DMoN–DPR (DVE) 9.74 71.30 43.98 47.36
383 DMoN–DPR (DVE) 12.58 72.50 42.83 45.78
277 DMoN–DPR (DVE) 13.81 69.51 40.61 45.21
274 DMoN–DPR (DVE) 10.97 73.71 46.42 47.92
188 DMoN–DPR (DVE) 9.64 73.99 51.33 56.97
796 DMoN–DPR (DVE) 10.80 71.35 42.74 43.29
Mean $\pm$ Std 10.93 $\pm$ 1.34 72.05 $\pm$ 1.39 44.37 $\pm$ 3.69 47.32 $\pm$ 4.63
E.2 CiteSeer
The following tables list the performance of DiffPool, MinCut pooling, DMoN and different DMoN–DPR configurations across different seeds on the CiteSeer dataset.
Table 18: Results obtained by using DiffPool pooling on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DiffPool 8.24 64.76 33.99 48.98
550 DiffPool 7.27 70.84 38.27 51.01
243 DiffPool 6.59 66.78 34.97 51.19
16 DiffPool 10.43 68.61 34.15 47.16
716 DiffPool 7.01 70.87 33.12 46.30
383 DiffPool 9.49 56.09 28.78 45.91
277 DiffPool 9.58 64.40 29.61 41.80
274 DiffPool 6.50 71.68 32.23 44.80
188 DiffPool 7.18 66.15 32.83 48.86
796 DiffPool 6.94 66.72 36.02 52.33
Mean $\pm$ Std 7.92 $\pm$ 1.42 66.69 $\pm$ 4.53 33.40 $\pm$ 2.82 47.83 $\pm$ 3.27
Table 19: Results obtained by using MinCut pooling on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 MinCut 5.45 76.59 33.17 45.03
550 MinCut 5.34 75.36 24.92 35.92
243 MinCut 5.01 76.67 30.26 42.82
16 MinCut 7.21 73.20 18.09 26.08
716 MinCut 7.01 74.82 22.11 30.09
383 MinCut 6.04 75.25 23.28 34.72
277 MinCut 7.05 73.86 17.40 25.14
274 MinCut 5.32 76.16 25.47 38.90
188 MinCut 6.44 75.23 32.94 42.54
796 MinCut 7.05 74.93 24.43 31.51
Mean $\pm$ Std 6.19 $\pm$ 0.86 75.21 $\pm$ 1.11 25.21 $\pm$ 5.52 35.28 $\pm$ 7.05
@@ E.2 CiteSeer
993 DMoN 6.81 74.92 20.78 26.09
550 DMoN 4.11 75.53 26.91 42.64
243 DMoN 4.28 75.57 32.12 45.74
16 DMoN 4.64 77.22 33.11 44.56
716 DMoN 4.53 76.24 36.95 50.12
383 DMoN 4.24 75.55 33.25 47.00
277 DMoN 4.26 75.63 29.25 42.33
274 DMoN 4.48 75.77 30.83 45.65
188 DMoN 6.17 72.74 24.52 36.01
796 DMoN 5.12 75.34 28.98 44.42
Mean $\pm$ Std 4.86 $\pm$ 0.91 75.45 $\pm$ 1.13 29.67 $\pm$ 4.70 42.46 $\pm$ 6.82
@@
993 DMoN–DPR (D) 7.18 74.27 23.90 27.73
550 DMoN–DPR (D) 4.22 75.80 26.23 41.98
243 DMoN–DPR (D) 3.82 76.02 35.91 50.13
16 DMoN–DPR (D) 4.83 76.31 32.78 44.47
716 DMoN–DPR (D) 4.57 76.13 36.99 50.14
383 DMoN–DPR (D) 4.81 75.97 33.02 45.96
277 DMoN–DPR (D) 3.87 75.75 29.37 42.34
274 DMoN–DPR (D) 4.15 76.06 30.88 45.48
188 DMoN–DPR (D) 6.37 73.98 25.00 36.47
796 DMoN–DPR (D) 5.71 75.26 26.83 39.93
Mean $\pm$ Std 4.95 $\pm$ 1.12 75.55 $\pm$ 0.81 30.09 $\pm$ 4.57 42.46 $\pm$ 6.69
@@
993 DMoN–DPR (V) 6.83 74.31 21.24 28.03
550 DMoN–DPR (V) 4.04 75.60 26.97 42.75
243 DMoN–DPR (V) 4.28 75.41 32.23 45.77
16 DMoN–DPR (V) 5.14 76.33 34.01 46.10
716 DMoN–DPR (V) 4.57 76.13 36.89 50.11
383 DMoN–DPR (V) 4.15 75.38 33.48 47.59
277 DMoN–DPR (V) 4.35 75.42 28.95 42.12
274 DMoN–DPR (V) 4.59 75.13 30.92 46.21
188 DMoN–DPR (V) 6.39 73.03 24.67 35.88
796 DMoN–DPR (V) 5.21 74.93 28.27 43.93
Mean $\pm$ Std 4.96 $\pm$ 0.96 75.17 $\pm$ 0.94 29.76 $\pm$ 4.71 42.85 $\pm$ 6.46
@@
993 DMoN–DPR (E) 7.27 73.95 21.36 28.28
550 DMoN–DPR (E) 4.15 75.67 26.79 42.59
243 DMoN–DPR (E) 4.31 75.39 32.27 45.75
16 DMoN–DPR (E) 4.92 76.51 34.13 46.04
716 DMoN–DPR (E) 4.48 76.00 37.03 50.46
383 DMoN–DPR (E) 4.46 75.36 32.82 46.38
277 DMoN–DPR (E) 4.20 75.58 29.78 43.08
274 DMoN–DPR (E) 4.26 75.81 31.63 46.59
188 DMoN–DPR (E) 6.22 73.30 25.10 36.09
796 DMoN–DPR (E) 5.07 74.98 28.59 44.38
Mean $\pm$ Std 4.93 $\pm$ 1.03 75.25 $\pm$ 0.96 29.95 $\pm$ 4.63 42.96 $\pm$ 6.35
@@
993 DMoN–DPR 7.54 73.39 23.91 29.07
550 DMoN–DPR 4.35 75.62 26.61 42.29
243 DMoN–DPR 3.84 75.78 36.17 50.51
16 DMoN–DPR 4.86 76.09 33.03 44.97
716 DMoN–DPR 4.55 76.03 38.13 51.55
383 DMoN–DPR 4.90 75.53 33.12 46.09
277 DMoN–DPR 4.06 75.33 29.89 42.79
274 DMoN–DPR 4.28 75.82 31.51 46.34
188 DMoN–DPR 6.50 73.71 25.38 36.63
796 DMoN–DPR 6.04 74.53 27.25 42.22
Mean $\pm$ Std 5.09 $\pm$ 1.21 75.18 $\pm$ 0.97 30.50 $\pm$ 4.72 43.25 $\pm$ 6.58
@@
993 DMoN–DPR (DE) 7.45 73.34 24.03 29.50
550 DMoN–DPR (DE) 4.37 75.63 26.92 42.46
243 DMoN–DPR (DE) 3.89 75.81 36.33 50.53
16 DMoN–DPR (DE) 5.07 75.80 32.09 43.42
716 DMoN–DPR (DE) 4.33 75.97 38.51 52.05
383 DMoN–DPR (DE) 4.83 75.58 32.59 45.72
277 DMoN–DPR (DE) 4.02 75.47 30.22 43.10
274 DMoN–DPR (DE) 4.35 75.80 31.50 46.10
188 DMoN–DPR (DE) 6.35 73.93 25.29 36.35
796 DMoN–DPR (DE) 6.06 74.49 27.23 42.08
Mean $\pm$ Std 5.07 $\pm$ 1.17 75.18 $\pm$ 0.92 30.47 $\pm$ 4.70 43.13 $\pm$ 6.53
@@
993 DMoN–DPR (VE) 7.51 73.51 21.85 28.66
550 DMoN–DPR (VE) 4.35 75.43 26.50 42.19
243 DMoN–DPR (VE) 4.48 75.15 32.58 46.11
16 DMoN–DPR (VE) 4.79 76.51 33.09 45.03
716 DMoN–DPR (VE) 4.50 75.95 37.46 51.15
383 DMoN–DPR (VE) 4.61 74.94 33.20 46.98
277 DMoN–DPR (VE) 4.24 75.33 30.08 43.40
274 DMoN–DPR (VE) 4.37 75.69 31.72 46.90
188 DMoN–DPR (VE) 6.39 72.92 25.40 36.39
796 DMoN–DPR (VE) 5.23 74.71 29.84 45.84
Mean $\pm$ Std 5.05 $\pm$ 1.07 75.01 $\pm$ 1.09 30.17 $\pm$ 4.53 43.27 $\pm$ 6.40
@@
993 DMoN–DPR 7.45 73.17 23.84 29.48
550 DMoN–DPR 4.15 75.58 26.66 42.48
243 DMoN–DPR 3.91 75.68 36.30 50.67
16 DMoN–DPR 5.23 75.44 31.58 42.92
716 DMoN–DPR 4.59 75.96 37.14 50.08
383 DMoN–DPR 4.75 75.43 32.95 45.93
277 DMoN–DPR 4.44 75.10 29.34 41.94
274 DMoN–DPR 4.35 75.91 31.30 46.10
188 DMoN–DPR 6.52 73.57 26.09 37.49
796 DMoN–DPR 6.24 74.21 27.32 42.91
Mean $\pm$ Std 5.16 $\pm$ 1.18 75.01 $\pm$ 1.00 30.25 $\pm$ 4.41 43.00 $\pm$ 6.16
E.3 PubMed
The following tables list the performance of MinCut pooling, DMoN and different DMoN–DPR configurations across different seeds on the PubMed dataset.
Table 28: Results obtained by using MinCut pooling on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 MinCut 12.82 53.38 23.17 41.29
550 MinCut 8.41 56.48 21.15 44.79
243 MinCut 12.52 53.77 23.64 40.72
16 MinCut 12.60 53.65 23.65 39.98
716 MinCut 10.43 55.36 21.44 37.44
383 MinCut 12.62 53.67 23.85 41.18
277 MinCut 8.32 56.56 20.47 44.08
274 MinCut 8.33 56.46 20.25 44.46
188 MinCut 12.52 53.76 23.34 38.89
796 MinCut 12.79 53.53 23.84 40.30
Mean $\pm$ Std 11.14 $\pm$ 2.04 54.66 $\pm$ 1.38 22.48 $\pm$ 1.47 41.31 $\pm$ 2.44